OpenStudy (anonymous):

If there are 6.02 x 1023 atoms per mole in any given substance, 4.70 x 1023 atoms of helium will represent how many moles of helium

OpenStudy (anonymous):

Division.

OpenStudy (anonymous):

Exactly. It's a conversion, and you know that you have \[X = 4.70 \times 10^{23} \ \text{atoms of He}\] and that \[\text{1 mole} = 6.022 \times 10^{23} \ \text{atoms}\] then the conversion is \[ \left( \frac{X \ \cancel{\text{atoms}}}{} \right)\left( \frac{\text{1 mole}}{6.022 \times 10^{23} \ \cancel{\text{atoms}} } \right) = \frac{X}{6.022 \times 10^{23}}\text{moles}=\text{# of moles}\]